OpenStudy (firejay5):

State the coordinates of the focus and the directrix equation for each parabola. *Show work and explain how you got what you got*

OpenStudy (firejay5):

\[y = \frac{ 1 }{ 8 }x^2 - 4\]

OpenStudy (tkhunny):

Firejay5! You been hiding? Personally, I don't like this form. Slightly more complicated than it has to be. \(y+4 = \dfrac{1}{8}(x-0)^{2}\) Now, THERE's a useful form. Vertex is staring at you. (0.-4) Also, the axis of symmetry. x = 0 Now, what about that Focus and Directrix?

OpenStudy (firejay5):

@satellite73 Can you work this problem out for me as an example, like the work for the focus and directrix and how you would get the answer